Topic: Implicit Differentiation


1. Find y’ in π‘₯3 + 2𝑦3 = 3π‘₯2𝑦.


2. Find the derivative of 𝑦 = βˆšπ‘ π‘–π‘›π‘₯𝑦

1
Expert's answer
2022-05-30T23:29:38-0400

1.


"x^3 + 2y^3 =3x^2y"

Differentiate both sides with respect to "x" and us the Chain Rule


"3x^2+6y^2y'=6xy+3x^2y'"

"x^2+2y^2y'=2xy+3x^2y'"

Solve for "y'"


"y'=\dfrac{x^2-2xy}{3x^2-2y^2}"


2.


"y =\sqrt{\sin(xy)}"

Differentiate both sides with respect to "x" and us the Chain Rule


"y'=\dfrac{1}{2\sqrt{\sin(xy)}}(\sin(xy))'"

"y'=\dfrac{\cos(xy)}{2\sqrt{\sin(xy)}}(xy)'"




"y'=\dfrac{\cos(xy)(y+xy')}{2\sqrt{\sin(xy)}}"

"2\sqrt{\sin(xy)}y'=y\cos(xy)+x\cos(xy)y'"

Solve for "y'"

"y'=\dfrac{y\cos(xy)}{2\sqrt{\sin(xy)}-x\cos(xy)}"